# Structure-Aware Contrastive Learning with Fine-Grained Binding Representations for Drug Discovery

### v1

@@ Abstract
I Introduction
@@
II Methodology
Figure 1: Model Architecture. Our framework encodes drug and protein sequences independently using frozen pre-trained models. The resulting embeddings are aggregated through an attention-based module and optimized via a contrastive learning objective. Cosine similarity between drug and protein representations is used to compute interaction probabilities, which support discriminative prediction of drug-target interactions and enable large-scale virtual screening.
II-A Representations of Drugs and Protein Targets
II-A1 Drug and protein encoders
As shown in Figure 1, the initial inputs of our model are string-based representations of drugs and proteins. For a protein sequence $P$, we adopt the structure-aware vocabulary [8], where each residue is replaced by one of 441 tokens that couple amino acids with 3-D geometric features. This design alleviates the lack of explicit structural information in amino acid sequences. For a drug sequence $D$, we employ SELFIES [7], a robust chemical string representation where every valid string representation can be unambiguously decoded into a chemically valid molecular graph, thereby avoiding invalid structures that may arise from SMILES.
Two frozen pre-trained encoders are employed to transform input sequences into representations. Protein sequences are encoded by Saport [8], which jointly models amino acid tokens and structural features to produce embeddings $\mathbf{P}$. Drug SELFIES sequences are processed by SELFormer [9], yielding embeddings $\mathbf{D}$. These embeddings capture cross-modal semantics relevant to molecular binding. The modular framework supports substitution with alternative language models, such as ESM-2 [10] for proteins or ChemBERTa-2 [11] for molecules.
II-A2 Attention-based pooling
@@
Formally, let $\mathbf{H}\in\mathbb{R}^{n\times d}$ denote the sequence of $n$ token embeddings with hidden size $d$. A learnable class token $\mathbf{q}\in\mathbb{R}^{1\times d}$ serves as the query against projected keys and values:
$\mathbf{K}=\mathbf{H}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{H}\mathbf{W}_{V},\qquad\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d\times d}.$ (1)
@@
$\bm{\alpha}=\mathrm{softmax}\!\left(\tfrac{\mathbf{q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\in\mathbb{R}^{1\times n},\qquad\tilde{\mathbf{z}}=\bm{\alpha}\mathbf{V}\in\mathbb{R}^{1\times d}.$ (2)
A residual connection with the class token, followed by a lightweight feed-forward network (FFN), produces the final representation:
$\mathbf{h}=\mathrm{FFN}(\tilde{\mathbf{z}}+\mathbf{q})\in\mathbb{R}^{1\times d}.$ (3)
@@
II-B Contrastive Learning for Drug-Target Interactions
We aligned drug and protein embeddings in a shared latent space via a symmetric contrastive objective. Given $N$ drug–target pairs, pooled embeddings $\mathbf{h}_{d}^{i}$ and $\mathbf{h}_{t}^{i}$ are projected to co-embeddings
@@ II-B Contrastive Learning for Drug-Target Interactions
$p(t=i\mid d=i)=\frac{\exp\!\left(\alpha\,\hat{\mathbf{Z}}_{d}^{i}\cdot\hat{\mathbf{Z}}_{t}^{i}\right)}{\sum_{j=1}^{N}\exp\!\left(\alpha\,\hat{\mathbf{Z}}_{d}^{i}\cdot\hat{\mathbf{Z}}_{t}^{j}\right)},$ (5)
with analogous formulation for target-to-drug. The overall loss is the symmetric InfoNCE:
@@
II-C Bilinear Attention for Interaction Representation
A bilinear attention network (BAN) captures token-level interactions between drug and target. Given token embeddings $\mathbf{H}_{d}\!\in\!\mathbb{R}^{n_{d}\times d}$ and $\mathbf{H}_{t}\!\in\!\mathbb{R}^{n_{t}\times d}$, we first obtain projected features $\mathbf{D}=\phi(\mathbf{H}_{d}\mathbf{U})$ and $\mathbf{T}=\phi(\mathbf{H}_{t}\mathbf{V})$ with $\mathbf{U},\mathbf{V}\!\in\!\mathbb{R}^{d\times r}$ and pointwise nonlinearity $\phi$.
$\mathbf{A}=\mathrm{softmax}\!\big(\mathbf{D}\mathbf{T}^{\top}\big)\in\mathbb{R}^{n_{d}\times n_{t}}$ (7)
$\mathbf{f}=\sum_{i=1}^{n_{d}}\Big(\mathbf{D}_{i}\odot(\mathbf{A}\mathbf{T})_{i}\Big)\in\mathbb{R}^{r}$ (8)
A single-head BAN uses (7)–(8), and multi-glimpse BAN repeats them with separate $(\mathbf{U}_{g},\mathbf{V}_{g})$ and concatenates $\mathbf{f}_{g}$. Finally, an MLP maps $\mathbf{f}$ to the interaction probability.
III Experiments
III-A Datasets and Baselines
The framework is assessed using DTI prediction and virtual screening. Following the DrugBAN [13] setting, BindingDB [5], BioSNAP [14], and Human [15, 16] datasets are employed for DTI prediction, where each dataset is split into training, validation, and test sets with a 7:1:2 ratio, and evaluated using 5-fold cross-validation ($k=5$). Virtual screening is conducted on the LIT-PCBA benchmark [17]. Dataset statistics are reported in Table I.
@@ III-A Datasets and Baselines
III-B Implementation Details
Training utilizes the AdamW optimizer with a learning rate of $5\times 10^{-5}$, weight decay of $1\times 10^{-4}$, and batch size of 64. Early stopping is applied after 20 epochs without improvement, with a maximum of 200 epochs. Drug and protein encoders generate embeddings of size 768 and 1280, respectively, which are linearly mapped to a shared 1024-dimensional latent space. Interaction modeling is performed using a bilinear attention module with 8 learnable heads. A dropout rate of 0.05 is applied throughout the network.
TABLE I: Experimental dataset statistics
@@ III-B Implementation Details
TABLE II: Performance comparison of DTI prediction methods on BindingDB, Human, BioSNAP datasets
@@
TABLE III: Performance comparison on LIT-PCBA benchmark
@@
IV Results & Discussion
IV-A Performance on DTI
Across three DTI datasets, the model achieves state-of-the-art results on Human and BioSNAP benchmarks. On BindingDB, performance aligns closely with DrugBAN and remains competitive with leading baselines. These results demonstrate consistent generalization and validate the effectiveness of the proposed architecture, as shown in Table II.
IV-B Performance on Virtual Screening
On the LIT-PCBA benchmark for virtual screening, the proposed method outperforms the previous state-of-the-art, DrugCLIP, as shown in Table III. SaBAN achieves an AUROC of 68.16%, yielding a 10.99% absolute improvement. For early recognition, SaBAN attains a BEDROC of 13.35%, exceeding DrugCLIP by more than twofold.
IV-C Ablation Study
To quantify the contribution of each module, we conducted ablation experiments by removing the learned aggregation layer (LA), the bilinear attention network (BAN), and the contrastive learning (CL) components. As shown in Figure 2, removing LA results in the largest performance degradation across all metrics. Excluding BAN or CL also leads to noticeable declines, confirming that both interaction modeling and representation regularization are important for robustness.
IV-D Visualization
We visualize the embedding distribution for the CYP3A4 target and its attention profile to interpret the representations. Figure 3(a) shows the T-SNE projection, where SaBAN produces embeddings with positive samples positioned closer to the binding pocket than DrugCLIP, indicating improved structural correspondence. Figure 3(b) displays the ligand-level attention weights, concentrating on functional residues within the pocket and providing evidence of model interpretability.
Figure 2: Comparison between SaBAN and its variants on BindingDB dataset
Figure 3: Visualization of feature distributions for the target CYP3A4 (a) using DrugCLIP and SaBAN models and the attention map (b) of its ligand
V Conclusion
We proposed a contrastive framework that combined cross-modal alignment with bilinear interaction modeling to extract drug–pocket relationships from tokenized sequences. The approach delivered state-of-the-art results on Human and BioSNAP datasets and maintained competitive performance on BindingDB. In virtual screening evaluations, it outperformed previous methods on LIT-PCBA, achieving notable improvements in AUROC and BEDROC. These capabilities supported efficient pre-screening using accessible one-dimensional sequences augmented with novel structure-aware representations.
VI ACKNOWLEDGEMENTS

### v2

@@ Abstract
Index Terms—  Drug discovery, Drug-target interactions, Attention network, Contrastive learning
1 Introduction
@@
2 Methodology
Fig. 1: Model Architecture. Our framework encodes drug and protein sequences independently using frozen pre-trained models. The resulting embeddings are aggregated through an attention-based module and optimized via a contrastive learning objective. Cosine similarity between drug and protein representations is used to compute interaction probabilities, which support discriminative prediction of drug-target interactions and enable large-scale virtual screening.
2.1 Representations of Drugs and Protein Targets
2.1.1 Drug and protein encoders
As shown in Fig. 1, the initial inputs of our model are string-based representations of drugs and proteins. For a protein sequence $P$, we adopt the structure-aware vocabulary [21], where each residue is replaced by one of 441 tokens that couple amino acids with 3-D geometric features. This design alleviates the lack of explicit structural information in amino acid sequences. For a drug sequence $D$, we employ SELFIES [11], a robust chemical string representation where every valid string representation can be unambiguously decoded into a chemically valid molecular graph, thereby avoiding invalid structures that may arise from SMILES.
Two frozen pre-trained encoders are employed to transform input sequences into representations. Protein sequences are encoded by SaProt [21], which jointly models amino acid tokens and structural features to produce sequence embeddings $\mathbf{H}_{t}$. Drug SELFIES sequences are processed by SELFormer [26], yielding sequence embeddings $\mathbf{H}_{d}$. The encoders remain frozen during training to leverage their learned structural priors efficiently, while only the downstream interaction modules and projection layers are optimized. Our selection of encoders is grounded in controlled ablation studies.
2.1.2 Attention-based pooling
@@
Formally, let $\mathbf{H}\in\mathbb{R}^{n\times d_{h}}$ denote the sequence of $n$ token embeddings (i.e., $\mathbf{H}_{d}$ or $\mathbf{H}_{t}$) with hidden size $d_{h}$. A learnable class token $\mathbf{q}\in\mathbb{R}^{1\times d_{h}}$ serves as the query against projected keys and values:
$\mathbf{K}=\mathbf{H}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{H}\mathbf{W}_{V},\qquad\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d_{h}\times d_{h}}.$ (1)
@@
$\bm{\alpha}=\mathrm{softmax}\!\left(\tfrac{\mathbf{q}\mathbf{K}^{\top}}{\sqrt{d_{h}}}\right)\in\mathbb{R}^{1\times n},\qquad\tilde{\mathbf{z}}=\bm{\alpha}\mathbf{V}\in\mathbb{R}^{1\times d_{h}}.$ (2)
A residual connection with the class token, followed by layer normalization and a lightweight feed-forward network (FFN) with dropout rate 0.05, produces the final representation:
$\mathbf{h}=\mathrm{FFN}(\mathrm{LayerNorm}(\tilde{\mathbf{z}}+\mathbf{q}))\in\mathbb{R}^{1\times d_{h}}.$ (3)
@@
2.2 Contrastive Learning for Drug-Target Interactions
We align drug and protein embeddings in a shared latent space via a symmetric contrastive objective. Given $N$ drug–target pairs, pooled embeddings $\mathbf{h}_{d}^{i}$ and $\mathbf{h}_{t}^{i}$ are projected to co-embeddings via learnable linear projections $C_{d}\colon\mathbb{R}^{d_{h}}\to\mathbb{R}^{d_{\mathrm{proj}}}$ and $C_{t}\colon\mathbb{R}^{d_{h}}\to\mathbb{R}^{d_{\mathrm{proj}}}$:
@@ 2.2 Contrastive Learning for Drug-Target Interactions
$p(t=i\mid d=i)=\frac{\exp\!\left(\tau\,\hat{\mathbf{Z}}_{d}^{i}\cdot\hat{\mathbf{Z}}_{t}^{i}\right)}{\sum_{j=1}^{N}\exp\!\left(\tau\,\hat{\mathbf{Z}}_{d}^{i}\cdot\hat{\mathbf{Z}}_{t}^{j}\right)},$ (5)
where $\tau$ is a temperature scaling parameter (set to 1 in our experiments). An analogous probability $p(d=i\mid t=i)$ is computed for target-to-drug alignment. The overall loss is the symmetric InfoNCE:
@@
2.3 Bilinear Attention for Interaction Representation
A bilinear attention network (BAN) captures token-level interactions between drug and target. Given token embeddings $\mathbf{H}_{d}\!\in\!\mathbb{R}^{n_{d}\times d_{h}}$ and $\mathbf{H}_{t}\!\in\!\mathbb{R}^{n_{t}\times d_{h}}$, we first obtain projected features $\mathbf{D}_{b}=\phi(\mathbf{H}_{d}\mathbf{W}_{U})$ and $\mathbf{T}_{b}=\phi(\mathbf{H}_{t}\mathbf{W}_{V})$ with learnable weights $\mathbf{W}_{U},\mathbf{W}_{V}\!\in\!\mathbb{R}^{d_{h}\times r}$ and pointwise nonlinearity $\phi$.
$\mathbf{A}=\mathrm{softmax}\!\big(\mathbf{D}_{b}\mathbf{T}_{b}^{\top}\big)\in\mathbb{R}^{n_{d}\times n_{t}}$ (7)
$\mathbf{f}=\sum_{i=1}^{n_{d}}\Big((\mathbf{D}_{b})_{i}\odot(\mathbf{A}\mathbf{T}_{b})_{i}\Big)\in\mathbb{R}^{r}$ (8)
where $\odot$ denotes element-wise multiplication.
A single-head BAN uses (7)–(8), and multi-glimpse BAN repeats them with separate $(\mathbf{W}_{U,g},\mathbf{W}_{V,g})$ and concatenates $\mathbf{f}_{g}$. Finally, an MLP maps $\mathbf{f}$ to the interaction probability.
3 Experiments
3.1 Datasets and Baselines
The framework is assessed using DTI prediction and virtual screening. Following the DrugBAN [1] setting, BindingDB [14], BioSNAP [31], and Human [13, 4] datasets are employed for DTI prediction, where each dataset is split into training, validation, and test sets with a 7:1:2 ratio, and evaluated using 5-fold cross-validation. Virtual screening is conducted on the LIT-PCBA benchmark [23]. Dataset statistics are reported in Table 1. Strict data separation is maintained to prevent leakage. For DTI tasks, we ensure training, validation, and test sets are disjoint at the pair level. For virtual screening, we follow the official split protocols based on target identity, ensuring unseen targets during evaluation.
@@ 3.1 Datasets and Baselines
3.2 Implementation Details
The model is optimized via AdamW with a learning rate of $5\times 10^{-5}$, a weight decay of $1\times 10^{-4}$, and a batch size of 64. To prevent overfitting, early stopping is triggered after a patience of 20 epochs, with a maximum training duration of 200 epochs. The drug and protein encoders generate embeddings of size 768 and 1280, respectively, which are subsequently projected into a shared 1024-dimensional latent space via linear transformation. For interaction modeling, a bilinear attention module with 8 heads is employed, alongside a global dropout rate of 0.05.
Table 1: Experimental dataset statistics
@@ 3.2 Implementation Details
Table 2: Performance comparison of DTI prediction methods on BindingDB, Human, BioSNAP datasets
@@
Table 3: Performance comparison on LIT-PCBA benchmark
@@
4 Results & Discussion
4.1 Performance on DTI
Across three DTI datasets, the model achieves state-of-the-art results on Human and BioSNAP benchmarks. On BindingDB, performance aligns closely with DrugBAN and remains competitive with leading baselines. These results demonstrate consistent generalization and validate the effectiveness of the proposed architecture, as shown in Table 2.
4.2 Performance on Virtual Screening
On the LIT-PCBA benchmark for virtual screening, the proposed method outperforms the previous state-of-the-art, DrugCLIP, as shown in Table 3. SaBAN achieves an AUROC of 68.16%, yielding a 10.99% absolute improvement. For early recognition, SaBAN attains a BEDROC of 13.35%, exceeding DrugCLIP by more than twofold.
4.3 Ablation Study
To quantify the contribution of each module, we conducted ablation experiments by removing the learned aggregation layer (LA), the bilinear attention network (BAN), and the contrastive learning (CL) components. As shown in Fig. 2, removing LA results in the largest performance degradation across all metrics. Excluding BAN or CL also leads to noticeable declines, confirming that both interaction modeling and representation regularization are important for robustness.
4.4 Visualization
We visualize the embedding distribution for the CYP3A4 target and its attention profile to interpret the representations. Fig. 3(a) shows the T-SNE projection, where SaBAN produces embeddings with positive samples positioned closer to the binding pocket than DrugCLIP, indicating improved structural correspondence. Fig. 3(b) displays the ligand-level attention weights, concentrating on functional residues within the pocket and providing evidence of model interpretability.
Fig. 2: Comparison between SaBAN and its variants on BindingDB dataset.
Fig. 3: Visualization of feature distributions for the target CYP3A4 (a) using DrugCLIP and SaBAN models and the attention map (b) of its ligand.
5 Conclusion
We propose a contrastive framework that combined cross-modal alignment with bilinear interaction modeling to extract drug–pocket relationships from tokenized sequences. The approach delivered state-of-the-art results on Human and BioSNAP datasets and maintained competitive performance on BindingDB. In virtual screening evaluations, it outperformed previous methods on LIT-PCBA, achieving notable improvements in AUROC and BEDROC. These capabilities supported efficient pre-screening using accessible one-dimensional sequences augmented with novel structure-aware representations.
6 ACKNOWLEDGEMENTS